what is the average friction force when a person has a usual 1km walk?

Grade:11

1 Answers

Sonu Sutradhar
5 years ago
As you have mentioned “USUAL 1 km WALK” so it should be kept in mind that the velocity of the person is constant.It is so because the person has a walk of 1 km very frequently or maybe regularly so his body muscles have attained a constant velocity for the path=1km through CONDITIONED REFLEX ACTION.
So it can be said that acceleration of the person(a)=dv/dt=0
We know,
               F=ma
               F=m*0=0
So the Average force of friction when a person has a usual walk of 1 km is ZERO